OpenStudy (anonymous):

Find the common difference of the arithmetic sequence. 25, 18, 11, 4, . . .

OpenStudy (anonymous):

help Please

OpenStudy (misty1212):

HI Subtract one term from the previous one

OpenStudy (anonymous):

i did that but i could not get the answer right away:(

rishavraj (rishavraj):

if a1, a2, a3, a4 ..... are in AP then common difference = a2 - a1 = a3 - a2